OpenStudy (anonymous):

Given the function f(x) = 2(x + 4), find x if f(x) = 20. 6 48 14 8

mathslover (mathslover):

\[\huge{f(x) =2(x+4)}\] \[\huge{2(x+4)=20}\] \[\huge{10=x+4}\] \[\huge{x=6}\]

OpenStudy (anonymous):

the answer is six Plug 20 in for F(X) and solve for "X" 20=2(x+4) 10=x+4 x=6